### v1

Abstract
The quality of Optical Music Recognition (OMR) systems is a rather difficult magnitude to measure. There is no lingua franca shared among OMR datasets that allows to compare systems’ performance on equal grounds, since most of them are specialised on certain approaches. As a result, most state-of-the-art works currently report metrics that cannot be compared directly. In this paper we identify the need of a common music representation language and propose the Music Tree Notation (MTN) format, thanks to which the definition of standard metrics is possible. This format represents music as a set of primitives that group together into higher-abstraction nodes, a compromise between the expression of fully graph-based and sequential notation formats. We have also developed a specific set of OMR metrics and a typeset score dataset as a proof of concept of this idea.
keywords:
Optical Music Recognition, Evaluation, Datasets, Computer Vision
1 Introduction
Written music has been part of the cultural heritage of humankind for many centuries. From neumes as old as a thousand of years to current-time western notation scores, people have built methods to make this most ephemeral of arts persistent through the passing of time. Rivers of ink have been imprinted on libraries worth of paper in order to preserve music along the ages, and much of it has indeed survived until today.
It is therefore unsurprising that, given the sheer volume of notably interesting (and oftentimes, forgotten) pieces of music that have been endowed to our current generations, scholars have turned their attention to computers to aid them in the endeavour of preserving and analysing them. The task that concerns us in this work is Optical Music Recognition (OMR), which is the first and therefore fundamental step for the computational analysis of written scores; converting images or scans of music into a defined format a computer can process Calvo-Zaragoza et al (2021).
Modern-day OMR has been greatly enriched with the advent of the powerful Deep Learning techniques developed during the last decade Calvo-Zaragoza et al (2021); Rebelo et al (2012). Nevertheless, the field itself remains fragmented, with few researchers fully devoted to it and each of them developing their own unique point of view and methodology Pacha (2018); Calvo-Zaragoza et al (2021). This is particularly evident when analysing the available datasets Pacha (2017), as can be seen in Table 1; most of them are restricted to specific steps or approaches Parada-Cabaleiro et al (2017); Tuggener et al (2018); Calvo-Zaragoza and Rizo (2018a) and almost none of them are compatible among each other (the most notable exception being the recent Shatri and Fazekas (2021), which incorporates its ground truth in multiple formats).
One of the main points of disagreement among OMR researchers is the matter of the evaluation of models Hajič and Pecina (2017); Hajič jr. (2018); Mengarelli et al (2019); Calvo-Zaragoza et al (2021), which is but the visible consequence of many long-standing issues – some related to the inherent difficulty of the field (for an example, see Figure 1), but others related to the diversity of plausible approaches to music recognition.
Figure 1: Fragment of a scan of a 1910 edition of Claude Debussy’s La Danse de Puck from the first volume of his Préludes. In this fragment, many of the reasons why modelling OMR scores is difficult can be seen – a two-staff system with voices that move along both of them, many small symbols concentrated in limited areas of the score, many independent voices sounding in unison, elements spanning multiple measures, chords, etc.
Evaluation of OMR models is currently performed on a per-methodology basis Calvo-Zaragoza et al (2021); Rebelo et al (2012). If one considers the goal of OMR to produce a computational representation of music – whichever that final representation might be  [^1] Byrd et al. Byrd and Isaacson (2003) collected a very extensive and detailed set of requirements for a music notation system that is suitable for both score engraving and OMR, but to the best of our knowledge there is no recognition system that formally or explicitly applies them. – one would expect the metrics to obey a criterion of either some aspect of music understanding or quantifying the correctness of the final representation. Instead, most metrics revolve around measuring the fidelity of intermediate representations devoid of musical meaning, disregarding the full reconstruction of music scores altogether. Moreover, the benchmarks this evaluation metrics are computed upon are rarely widespread, with each methodology using their own. Commercial OMR systems have also been very vague about the specifics of their capacity.
With the aim of setting a step forward towards unifying the efforts of the community, we propose a way to set a common framework for OMR. We believe the first step is to set a target final representation that accommodates a significant amount of use cases within Common Western Music Notation (CWMN). We explicitly avoid trying to cover the entirety of OMR, as the nature of notated music is diverse and heavily dependent on the time it was written and the instrument it was written for. Once this target representation is chosen, a way of easily working with it has to be developed. Currently used engraving formats are quite exhaustive, verbose and option-filled, which allows expressing the same score in a variety of ways – something that makes them quite impractical for direct comparison. Thus, we propose an intermediate representation that translates directly to the desired output but facilitates OMR operations – including analysis, evaluation and composition of results – regardless of the underlying approach. We propose doing so through a notation based on the idea of building a tree structure; leaves are the objects in the music score and intermediate nodes are their composition. The idea is exploiting the expression advantages of graph-based approaches while simplifying the kinds of relationships that such system can model. Incorporating a common notation format is also a way of unifying all existing datasets so that they can be both combined and compared on equal grounds, regardless of the methodology they were originally designed to work with.
The contribution of this work can be summarised by the following claims:
  • We try to bridge the gap between the different benchmark suites in OMR literature with a universal tree-based notation format designed to represent musical scores at the graphical level.
  • We also present a common evaluation toolkit which aims towards unify existing benchmark OMR tasks for fairer comparison.
  • We have produced a typeset dataset  [^2] The data shall be published upon acceptance alongside the conversion and evaluation code. using public domain works with permissive licenses.
This paper is structured as follows. In section 2 some definitions for recurrent topics in the article are presented. Then, section 3 justifies the necessity for an OMR evaluation framework and debates about how that framework should be structured. In section 4, our proposed notation format is presented and described in detail, whereas in section 5 we describe the evaluation metrics to be employed with it. In section 6 we introduce the proof-of-concept dataset on typeset scores. Finally, section 7 closes the paper with some final remarks and future work.
2 Terminology
In this section we shall address some specific recurrent terms for the sake of clarity and cohesiveness of the text. Throughout this paper we shall adhere to the nomenclature and definitions found in Calvo-Zaragoza et al.’s Calvo-Zaragoza et al (2021).
By structured output we refer to an output encoding format that allows the replication of the input score holistically and unambiguously. This is in contrast to partial outputs such as MIDI, which only encodes note pitch, duration and velocity and omits any other indication present on the score.
Another relevant point is the definition of an end-to-end OMR model. Modern day literature employs this term frequently, but it must be noted that it is used with a slightly different meaning than the term itself actually conveys. The most widely recognised definition of the OMR pipeline is perhaps that of the review penned by Rebelo et al. Rebelo et al (2012), which states four distinct steps:
  1. 1.
    Image preprocessing.
  2. 2.
    Recognition of Musical Symbols.
  3. 3.
    Reconstruction of Musical Information.
  4. 4.
    Construction of a Musical Notation model.
Most end-to-end approaches do not actually perform all four steps; rather, these models stop at step 3 with a notation that encodes most of the score’s information directly or indirectly, but cannot be directly engraved (sometimes requiring some additional interpretation before all music semantics can be successfully produced). Whenever we address end-to-end approaches, we refer to these. In this work we mostly address offline use cases of OMR – that is, recognition of raster images of scores. Nevertheless, we believe some of our ideas are equally applicable in online use cases Calvo-Zaragoza and Oncina (2014).
3 The Necessity for a Common OMR Evaluation Framework
In this section we provide thorough arguments as to why having a comprehensive and common OMR evaluation system is necessary, while also providing some insights into what we deem is the preferable approach given the current circumstances of the field.
3.1 OMR Needs a Standardised Output
Most OMR research today does not tackle the problem of converting a model’s output to a symbolic representation that can be engraved into a proper score. In their review of the field in 2021, Calvo-Zaragoza et al. Calvo-Zaragoza et al (2021) go as far as claiming that no research OMR system that they know of is actually capable of performing such a conversion [^3] There are however commercial OMR systemsBitteur (2004) that can perform this step., and since its publication we have only acknowledged one system that can Shishido et al (2021). One of their main justifications for this is the fact that most structured music engraving formats are either not mature enough to cover all possible use cases or simply impractical for OMR tasks. In practice, what this means is that very few OMR systems spawned in research contexts are employed by final users. Furthermore, it hinders OMR systems’ evaluation potential, as omitting music semantics inevitably reduces the level of interpretability of the models’ outputs.
Going back to the matter of whether notation systems are practical or mature enough for recognition, there are a few points that need to be made. If the final goal of recognition is a structured encoding that attempts to replicate the original score with the highest fidelity, there are currently two realistic choices: MusicXML Good (2001) and MEI Roland (2002). A point can be made about LilyPond’s The LilyPond Developement Team (2014) ability to engrave scores from text written in a fairly succinct and descriptive fashion, but since the specification of the format is subject to change and the fact that it under-specifies some aspects of the score for the sake of usability, we do not think it is realistic to use it as a comprehensive final output. Some effort has also been devoted into extending Humdrum’s **kern notation to make it capable of engraving scores uniquely Ríos-Vila et al (2023), but some of the same limitations as those seen for Lilypond apply.
MusicXML and MEI are somewhat similar in their base design ideas, with the former being more widely used within music score repository platforms and engraving software and the latter in libraries and academic-oriented settings. Perhaps the most obvious difference between the two is the fact that MusicXML is designed with the idea of enabling MIDI playback out-of-the-box at the cost of some redundancy, whereas MEI is more focused on musicological and archival applications. At the end of the day, they both can model the most widely used subset of CWMN music and they both have limitations derived from their use of XML and CWMN’s inherent flexibility; it is difficult to express co-depending symbols as a tree-like hierarchy. Therefore, both choices are probably equal in value in terms of expressiveness.
Thus, we believe it to be a net positive to make all OMR applications interoperable with these formats by design. The first advantage is making it possible to export scores back and forth and therefore increasing the amount of data available for recognition while making OMR more appealing for the end-user. The second one is that the set of requirements of an OMR system is grounded to those of the target application, making it simpler to cooperate among the community as everybody is talking the same language. Finally, and thanks to the fact that the targeted output is well defined, unambiguous metrics for end-to-end score recognition can be defined.
The big question is then how to make it practical for OMR practitioners to target these (rather complex) formats in a streamlined fashion. Our belief is that it is best not to use them directly, but rather incorporating a simplified intermediate format that makes it easy to convert to the structured outputs and also sets a semantic frame that adjusts well to the necessities of OMR.
3.2 On Computational Music Representations and their Tradeoffs
The most flexible – and perhaps the most logical – computational representation of written music is most likely a graph, since it allows expressing $n$-dimensional arbitrary relationships between primitives and symbols. There exists a very widely known notation format based on this idea called Music Notation Graph (MuNG) introduced for the MUSCIMA++ Hajič and Pecina (2017). Some works have indeed succeeded in being able to reconstruct music structure from these graphs, such as Baró et al.’s Musigraph Baró et al (2022) or Garrido-Muñoz et al.’s Convolutional Recurrent Neural Networks Garrido-Munoz et al (2022). Nevertheless, to the best of our knowledge, there is no straightforward way of converting this graph output back into a structured encoding. Moreover, the cost of evaluation on graphs is high due to the complexity of matching graphs together.
At the other end of the spectrum lies the sequential music model, currently employed in some form or another by most end-to-end OMR approaches Calvo-Zaragoza et al (2017b); Calvo-Zaragoza and Rizo (2018b); Baró et al (2019); Alfaro-Contreras and Valero-Mas (2021); Ríos-Vila et al (2022b); Baró et al (2020); Torras et al (2021); Wen and Zhu (2022). These methods impose a relationship between the horizontal axis of the image input and the playback time of the score. However, this forces the insertion of music semantics in order to properly model specific musical concepts such as unison, as the only reliable way of aligning individual voices to their playback positions is by providing their starting times. This is, in fact, how ubiquitous formats such as MusicXML deal with this ambiguity – a “backup” token specifies that the internal clock of the engraver must be brought back a number of time partitions and start producing notes there. Current end-to-end OMR efforts are mostly restricted to either monophonic or homophonic music scores as a result.
This dichotomy of approaches defines a search space that is still being explored Alfaro-Contreras and Valero-Mas (2021); Alfaro-Contreras et al (2022); Baró et al (2019); Hajič and Pecina (2017). Our proposal is to try and exploit the advantages of both ends: the structural expresiveness of graphs and the ease of use of sequences. We propose modelling music as a tree structure, in which small semantic units (primitives) can be composed into complex high-level constructs (note groups, measures). The tree restriction makes it simpler to operate with, as the number of relationships is drastically reduced, relationship semantics are strictly hierarchical (and thus simpler to model) and its resemblance to the targeted formats makes conversion a simpler task. Also, imposing an ordering restriction within the tree as in a sequence makes matching computationally less demanding and allows modelling the idea of time and position somewhat implicitly.
3.3 The Complicated Endeavour of Music and Data
Table 1: Description of annotation formats available for OMR datasets designed for final output production (i.e. they are designed to reconstruct a full score in some form). Columns from left to right: the publication that spawned the dataset, the name it is usually referred to as, what type of music samples it contains and whether it contains bounding box annotations, pixel-mask annotations, notation graph (MuNG) among symbols, string-based output annotations and a final structured output (MusicXML or MEI).
Publication Dataset Name Dataset Type Boxes? Masks? Graph? String? Struct?
Hajič and Pecina (2017)
MUSCIMA++
Handwritten Contemporary
Tuggener et al (2018, 2020)
DeepScores
Typeset
Tuggener et al (2023)
RealScores
Scanned
Shatri and Fazekas (2021)
DoReMi
Typeset
Parada-Cabaleiro et al (2017)
SEILS
Typeset Scanned Mensural
Baró et al (2020)
Baró Synthetic
Typeset
Baró et al (2020)
Pau Llinás
Handwritten Historical
Calvo-Zaragoza and Rizo (2018b)
PRIMuS
Typeset
Calvo-Zaragoza and Rizo (2018a)
Camera PRIMuS
Typeset
One of the longest-standing and most crippling problems OMR has faced in the recent years is the lack of reliably annotated data for contexts other than recognition of typeset scores. Even in that case, the datasets are heavily opinionated and rarely incorporate structured music sources (see Table 1 for more detail).
In particular, there is a very noticeable lack of annotated data in CWMN set within historical contexts – that is, scores written from the 16th century up to the 20th. As can be seen in Table 1, only Baró et al.’s Baró et al (2020) Pau Llinás dataset is dedicated to music from this period. Ironically, some of the composers within it are among the most extensively studied musicians of all time: Bach, Beethoven, Mozart, Händel or Haydn, to name a few.
Given their importance to the art and craft of music, many efforts have been made to preserve their work. Some libraries have extensive repositories of historical music documents from a very broad period of time, some of which have devoted efforts specifically aimed at democratising access to big names in music noa (2023, 2008). Some of these original manuscripts are currently part of the public domain, which also enables them for use within OMR research.
The main question however is how to best incorporate this data into the field, as most of the original manuscripts do not have an associated ground truth that is useful for OMR. The most direct approach would be to collect a significant corpus of works and manually annotate it comprehensively (per-pixel symbol annotation, music score transcription, etc.). Nevertheless, this annotation is both extremely costly and time-consuming, and it entails an implicit choice of domain; scores that do not abide to the statistical properties of the collected corpus will still be intractable, and thus the manual annotation cycle would begin anew.
Therefore, our belief is that the way forward is not just about producing new hand-annotated datasets, but rather trying to exploit every single bit of data already publicly available. More specifically, we propose the creation of new OMR datasets based on existing transcriptions of public domain works. The goals are twofold: 1) to build a corpus of real scores soft-aligned (that is, a dataset where the position of measures within the page is known but the position of each individual object need not be) with existing transcriptions and 2) to enable the generation of old-looking synthetic scores, an avenue that is already being explored by Mayer et al. Mayer and Pecina (2021); Mayer (2022).
The downside of this approach is that producing symbol-level masks becomes considerably more complicated, which in practical terms shuts down the use of object detection-based methods in the short term. We nevertheless believe that this is probably going to be a temporary situation, as enabling recognition of these scores might potentially open the door for robust implicit segmentation or alignment systems with which an object-level dataset can be bootstrapped.
When it comes to the problem of alignment of scores, some issues do arise in the form of conflicting transcriptions and notation adaptations by editors or scholars. Nevertheless, heuristics can be set in place to detect consistent mistakes or usual deliberate changes – e.g. changes in the time signature, changes in the way voices are arranged on the staff, changes in clef and so on – and then apply the appropriate modifications back on the ground truth transcription. Those cases where the transcription is wrong are assumed to be minimal but possible nonetheless, as with any other human-produced ground truth.
Other roads for the use of non-annotated data currently exist to improve recognition efforts. In line with recent trends in the literature about building very large language models, many methods employ unlabelled input samples as some form of proxy task target. Examples of such techniques in Computer Vision related tasks include SimCLR Chen et al (2020) or Masked Autoencoders He et al (2022), which to the best of our knowledge have never been used in the context of OMR.
3.4 Evaluation of OMR Models
Having discussed the benefits of having a common notation format and the importance of producing a structured output, we now delve into the specifics of evaluation of OMR systems.
Unfortunately, most metrics used by the OMR community in modern day are fairly domain-specific and rarely take into account music semantics, a direct consequence of not targeting a musically-meaningful output format Mengarelli et al (2019). Authors that address OMR as an object detection problem report object detection metrics, whereas authors that address OMR as an image-to-sequence task rely on sequence edit distances. Consequently, there is no realistic way of comparing them since the semantics of the mistakes of the system are lost and the underlying notations are different. The few authors that do report musically-aware metrics restrict themselves to per-note pitch and/or time accuracy Baró et al (2019); Huang et al (2019).
Moreover, end users for OMR systems are not concerned so much about the absolute number of mistakes, but rather how costly or realistic it is to fix them, something that is hard to convey with metrics that treat all errors equally. A simple example of why this distinction is relevant is the following: imagine there are two available OMR systems, one that is unreliable providing correct note pitches but is otherwise very robust and another one which is very good at determining music semantics but tends to miss structural primitives. Instead of devoting efforts improving both systems independently, the outputs of both models can be combined and weighted according to their strengths to produce better results at essentially zero additional cost. This is something that has already been done for other similar tasks such as Handwritten Text Recognition Bertolami and Bunke (2008).
In this direction, there is also debate regarding the specific aspects of music that should be measured. Hajić jr. Hajič jr. (2018) argues that the best metric is one that summarises the entire notation system as one single value, reminiscent of the Symbol Error Rate (SER) metric for strings. We believe that, while such a metric is a good indicator of overall performance (and is definitely useful for benchmarking), it must not be considered on its own, as definite conclusions about the underlying models are very hard to make. Our belief is that there should be an array of metrics that are oriented at different aspects of the score that should be logged together, increasing the coverage of the system.
3.5 The Ideal Scope for Recognition
In this final section we discuss the effect of tackling OMR at different levels of detail.
At the current day there is no pre-defined scope to tackle the problem of OMR. Therefore researchers can choose whether to implement recognition at measure Baró et al (2020, 2022), line Calvo-Zaragoza and Rizo (2018a); Baró et al (2019); Wen and Zhu (2022) or page level Pacha et al (2018); Ríos-Vila et al (2022a); Zhang et al (2023). While there are practical advantages on each perspective, there is an arguing point that should be brought to discussion given the current situation of OMR and Computer Vision as a whole.
Current state-of-the-art object detection and recognition models struggle with both small Yu et al (2020); Pacha (2021); Baró et al (2022) and scale-variant symbols Zhang et al (2018). Music scores are quintessential examples of both phenomena: large symbols (slurs, ties, staves) coexist with very small ones (dots, fingering numbers), and writers draw symbols at their preferred relative scale without affecting their meaning – something that is particularly true for noteheads and stems – while also writing semantically-distinct symbols that change based on their size – grace notes. Moreover, the number of objects present in the input image also affects the recognition performance, a recurrent matter for a significant amount of pieces.
Since addressing visual recognition tasks at smaller scopes helps to obtain better results, we believe that OMR should first be approached at measure level, since this is the smallest semantic unit of music. The first immediate consequence is that layout analysis of music scores becomes a lot more important, as all successive recognition steps depend on finding the measures reliably on the page. Nevertheless, we argue that it is objectively a simpler problem than fully-fledged recognition: staves and bar lines are very prominent, very ubiquitous and can easily be identified heuristically Fujinaga (2004); Cardoso et al (2009); Fornés et al (2013); Bui et al (2014); Calvo-Zaragoza et al (2017a); Egozy and Clester (2022), as opposed to many of the other musical symbols. Therefore, by composing small semantic units and propagating shared semantics (key, time signature, clef) whenever needed, a full notation reconstruction can be obtained quite simply, as measures are mostly independent.
Tangentially, this idea also helps in the recognition by providing higher dpi images which are more informative, as there is a substantial part of the score that is usually plain white space. Some tricks have been developed to try to minimise these gaps, such as removing all horizontal space between primitives Zhang et al (2023). This is not something particularly useful with scores with noisy backgrounds (e.g. scans from old documents with signs of paper degradation, a recurrent matter within old music scores), but it has been proven to improve results substantially.
In this section we have provided the core ideas for a common OMR notation format and evaluation procedure. This shall be the logical foundation upon which our proposed solution, described in the following section, is built.
4 The Music Tree Notation Format
As stated in Introduction and in concordance to the necessities discussed in Section 3, we have designed a notation format that
  • enables comparison of diverse OMR methods on equal grounds,
  • normalises the set of music primitives to be recognised,
  • simplifies conversion to a final structured format and
  • facilitates the usage of non-OMR-specific data.
The base logic for the notation system is introduced in 4.1. The notation system is described in detail in 4.2.
Figure 2: Example showing a fragment of a measure in which the annotation format for attributes and staff-modifying elements is shown. Rectangular nodes represent primitives as tokens and rounded nodes are abstract elements.
4.1 Rationale
The basis for our approach to OMR evaluation and (optionally) end-to-end inference is the Music Tree Notation format (MTN). We believe the final goal of structured OMR is the reconstruction of the score at the visual domain. The core idea of this format is therefore to build a notation that exclusively models relationships between graphical symbols and defers inference of music semantics until a later stage. Only those high-level music concepts that are strictly required to reconstruct the score unambiguously are kept if and only if there is a direct graphical cue that allows straightforward inference.
Evaluation should be possible regardless of the scope of recognition of the evaluated system – page, line or measure. Nevertheless, a minimal semantic unit is still required in order to reconstruct music semantics afterwards and facilitate alignment of predictions and their groundtruth. Exploiting the assumption evaluation is performed on the visual domain and therefore application of semantic-altering objects is irrelevant, we propose using the measure as the minimal unit of music on which to evaluate.
MTN should act as a final target for OMR, through which to convert the score into any other high-level notation system. This is in order to facilitate OMR systems from research being deployed into real use-cases as well as enabling any source of structured data to be used in OMR applications easily. Therefore, MTN draws heavy inspiration from both MEI and MusicXML, but “normalises” their structure in such a way that every score can only be expressed in one way. This also allows implementation many of the already-studied evaluation requirement ideas from other authors, particularly Hajič jr. Hajič jr. et al (2016) and Byrd Byrd and Simonsen (2015).
Finally, music notation is very rich and tends to bend its own rules quite heavily. The goal for this notation and evaluation system is to accommodate a most significant subset of CWMN, but covering every single exception proves to be a daunting task. Byrd Byrd and Simonsen (2015) presents some specific cases such as a spanning slur with seven inflection points that is rather complicated to model in the logical domain. Therefore, our rationale is that support for these should not be prioritary. We believe it is better to have some solid tools on a core subset of music rather than trying to cover every possible case with highly complex and fragile ones as this makes adoption of the format much more difficult.
4.2 Description of the Format
A graphical representation of a simple measure engraved in MTN can be seen in Figure 2. The core element of this format is the Musical Primitive, a concept that is quite widespread in the OMR literature Baró et al (2016, 2020); Calvo-Zaragoza and Rizo (2018a); Tuggener et al (2018) and can be defined as any of the independent structural elements that may or may not be combined together to form a semantic unit in the music score. The set of musical primitives includes all graphical elements in a score that are self-contained and require no other symbols to convey meaning (this includes rests, clefs or time signatures), the set of graphical elements that compose notes (noteheads, stems, flags, dots, etc.) and other miscellaneous elements such as numbers for compound time signatures. Every primitive is given a unique work-level identifier.
These primitives associate together to form more abstract constructs. This is modelled in MTN using a tree-like structure of higher-order elements resembling an Abstract Syntax Tree (AST), which defines the set of dependencies among objects in the score. This idea emulates parsing the contents of the score using a grammar, enabling the bulk of tools and research on parsers, parser generators and AST analysis and processing to be used in the context of music. Furthermore, it is a structure that can be modelled very easily using widespread exchange formats such as XML or JSON. The structure of the format may be found in Table 2.
Table 2: Definition of elements within MTN and the overall structure of the format. Each of the lowest levels of the notation contain the various final tokens. For more information, refer to the implementation of the format.
Top Level Level 1 Level 2 Definition
Attributes
Any semantic change the score goes through.
Staff
Group of changes occuring on the same staff.
Clef
A change of clef in the score.
TimeSig
A change of time signature in the score.
Key
A change of key in the score.
Barline
A measure separator and its attached modifiers.
Direction
A single playback directive.
Note Group
Group of notes tied together by at leas one beam.
Chord
Set of notes in unison attached to the same stem.
Stem
Stem direction information and beams or flags attached to it.
Note
A notehead and its attached modifiers.
Rest
Rest token and its attached modifiers.
There are, however, some elements in music that break the tree-like structure assumption. These are elements that connect multiple notes together outside their local note group structure: slurs, ties, parentheses and tuplets, among others. Both MEI and MusicXML acknowledge this limitation and circumvent it through the use of identifiers. MTN is no different: it provides a unique starting and ending token for each side of the object and gives both ends the same identifier.
In order to emplace musical elements in their corresponding positions, MTN uses two magnitudes. Firstly, for every token a tuple of two integers denotes the staff the element belongs to and its position within the staff. The position is denoted counting the number of steps from the first ledger line below a staff. For those elements without a specific position (such as rests or stems), a null value is used. Secondly, for any object immediately below the class measure, an exact timing value is provided. It is measured in fractions of a quarter note from the start of the measure itself. This information is also provided for every chord in a note group even if this information can be inferred for the sake of simplifying evaluation procedures.
Figure 3: The first few measures on Claude Debussy’s first Arabesque. The single voice moves along both staves and is played using both hands on a piano. If one was to separate both staves, it would be impossible to extract the exact timing where each group of notes is to be played. Another interesting aspect of this piece is that many editions use implicit triplets, something that only can be inferred contextually by the fact that eighth notes are on groups of three.
This musical information is necessary for two reasons. The only visual cue for aligning elements in a polyphonic setup is that, at some point, addicional voices have at least one object that is perfectly aligned with one in the main voice. This is very hard to model cleanly with anything other than a graph. Moreover, the main voice of a measure may arbitrarily jump from staff to staff in a system, making timing generally impossible to resolve on a per-staff basis, as can be seen in Figure 3. A small trick to optimise sequence-based notation systems is that one only needs timing information for the first element in a group, as all successive chords’ timings can be inferred from the note types and tuplets attached to them (or their surrounding elements).
Doubly stemmed notes (such as in the last measure of the Arabesques example in Figure 3) can be interpreted like two notes playing at the same time. Therefore, they can be modelled using two distinct groups without needing to break the tree-like structure.
Finally, to enable unambiguous comparison, a reading order of sorts must be established. We propose the following ordering criterion:
  • By starting time counting from the beginning of the measure.
  • By top level class:
    1. 1.
      Attributes
    2. 2.
      Directions
    3. 3.
      Rests
    4. 4.
      Note Groups
    5. 5.
      Barlines
  • By staff position: first objects on upper staves and lower positions within them.
  • In case of Note Groups, by direction of the first stem: first stems looking upwards.
  • For anything else, token alphabetical order. This also guarantees stability of the notation if new token types are added.
4.3 Extensions
The main goal for the notation is to address symbolic-level OMR. Nevertheless, many practitioners use pipelined architectures where some partial results need to be annotated (most notably, object bounding boxes). Providing these annotations with MTN is straightforward: it is a matter of adding an “annotations” field with each token identifier linked to the desired information. This also includes some other relevant semantic cues such as voices, which we deem unnecessary for the task of score reconstruction but might be useful for some other music information retrieval tasks. A concern surrounding voice information is that its definition around music engraving software is more akin to a layer in a graphic design suite; it is used more as a hacky way of circumventing engraving limitations, making voice information in files like MusicXML rather arbitrary.
The presence of text in scores can also be treated separately from the rest of music semantics. Since its interaction with the score is very situational and depends a lot on the contents of the text itself, it is very difficult to devise a set of relationships that might be useful for most scores. Therefore, the only text that is to be represented in MTN are dynamic directives, which are common and well-defined. For the rest of the text such as tempo indications or other textual remarks, we believe it is best to append a set of text boxes in a similar way as bounding boxes are provided.
The format itself is still maturing phase and open to feedback from the community. We are open to extending it to accommodate the needs of OMR practitioners as they put it into test.
5 Evaluation Metrics
We propose a set of evaluation metrics that both acknowledge the existence of multiple paradigms for OMR while also setting ways to compare any structured output equally. These metrics draw inspiration from the currently used Symbol Error Rate and ideas from Hajič jr. Hajič jr. et al (2016). We have divided our proposed metrics as tiers depending on the abstraction level they address.
5.1 Tier 0: Methodology-Specific Metrics
The first tier is that devoted to specific pipeline metrics before performing any kind of reconstruction into the MTN format. The idea behind this tier is that comparison of methodology-specific metrics is still useful, since it might help identify strengths or shortcomings on certain applications and help quantify the effectiveness of methods converting into MTN format (which depending on the OMR methodology might be more or less nuanced).
Since the approaches to OMR are both diverse and unique, it does not make much sense to overly specify what metrics to employ under this tier, but rather use those metrics that are already being employed. Some examples of the kinds of metrics that should be reported here include (but are not restricted to):
  • For object detection methods, metrics such as mean average precision, F1 score, average intersection over union and so on.
  • For end-to-end string-based methods, metrics such as symbol error rate or string parseability (whether or not the output string can be interpreted correctly with whichever grammar describes its underlying language).
5.2 Tier 1: Primitive Level
The first set of metrics addresses the presence or absence of terminals within the MTN string. The main goal is determining the capacity of models of detecting objects and analysing token co-ocurrence phenomena. These metrics do not take into account structural matters, making them quick to compute. Given a set of unique terminals $T$, which is a subset of the full MTN vocabulary $V$, the multiset of predicted terminals within a sample $P=\left\{p_{1}\dots p_{n}:p_{i}\in T\right\}$ as the multiset of all terminals produced by a model and the multiset of ground truth terminals within a sample $G=\left\{g_{1}\dots g_{m}:g_{i}\in T\right\}$ as the multiset of all terminals existing in the groundtruth (both in no particular order), we define the following metrics:
  • Primitive-level precision
    $precision=\frac{\|P\cap G\|}{\|P\|}$ (1)
  • Primitive-level recall
    $recall=\frac{\|P\cap G\|}{\|G\|}$ (2)
Since the underlying datasets are quite unbalanced due to the presence of many dominating tokens (e.g. noteheads), these metrics are computed per-class for the entire dataset. In order to produce a single precision and recall measure, results are aggregated per-class using a weighted average, where the weights are the relative frequency of each token in the ground truth.
5.3 Tier 2: Structural Level
This level takes into account the structure of the produced MTN and compares it directly with that of the ground truth. A matching from ground truth elements to those present in the prediction is performed using a tree edit distance algorithm. In particular, since there is a restriction on the ordering of sibling labels, the $O\left(n^{3}\right)$ solution from Zhang and Sasha can be employed Zhang and Shasha (1989).
Given the following operations:
  • Substitution: Changing the label of a single node within the tree.
  • Deletion: Removing a single node of the tree and setting its children as siblings.
  • Insertion: Adding a new node under a parent one and setting a consecutive subsequence of its siblings as children.
Given a predicted tree and a ground truth tree whose set of vertices is $G$ and assuming an equal edit cost of 1 for all operations, the Tree Error Rate (TER) is defined as
$TER=\frac{S+D+I}{\|G\|}$ (3)
where $S$, $D$ and $I$ are the number of substitution, deletion and insertion operations required to produce the ground truth tree from the predicted tree. This metric is designed mostly for benchmarking and is defined by analogy to the ubiquitous Symbol Error Rate (SER).
5.4 Tier 3: Semantic Level
This level considers some aspects of music semantics, particularly those concerned with pitch and time. This tier depends on the matching extracted from the structural level in order to identify the association between objects and their ground truth correspondences. Since evaluation has to be possible at the measure level, we use a graphical definition of pitch, which is the position of the notehead within the system.
A note $n$ is defined as a 3-tuple $(p,t,d)\in\mathbb{Z}^{2}\times\mathbb{Q}\times\mathbb{Q}$, where $p$ is a tuple of the staff the note belongs to and its position within the staff, $t$ is the number of beats from the start of the measure and $d$ is the duration of the note in number of beats. Rests are considered notes whose pitch is irrelevant (in practical terms, they can be thought of having a unique shared number for pitch). Given the set of predicted notes $P$ and the set of ground truth notes $G$, the matching between $P$ and $G$ is defined as a set of tuples $M=(n_{p},n_{g})\in P\times G$ such that at most one tuple with either a specific $n_{p}$ or $n_{g}$ exists.
The Missing Note Rate (MNR) is defined as the ratio of ground truth notes that do not have a corresponding prediction. More formally,
$MNR=\frac{\|\left\{n_{g}\in G:(n_{p},n_{g})\notin M,\forall n_{p}\in P\right\}\|}{\|G\|}.$ (4)
Similarly, the False Positive Rate (FPR) is defined as the ratio of predicted notes that do not have a corresponding ground truth note. It is defined by analogy to MNR.
Pitch Precision (PP) is defined as the number of correctly predicted pitches w.r.t. the ground truth as defined by the matching $M$. Therefore,
$PP=\frac{\|\left\{(n_{p},n_{g})\in M:p_{n_{p}}=p_{n_{g}}\right\}\|}{\|M\|}.$ (5)
Time Precision (TP) and Duration Precision (DP) are defined analogously.
The Average Pitch Shift (APS) is defined as the average offset in pitch from the predicted note w.r.t. its corresponding ground truth note as defined by the matching $M$. Therefore,
$APS=\frac{1}{\|M\|}\sum_{\forall(n_{p},n_{g})\in M}p_{n_{p}}-p_{n_{g}}.$ (6)
Time Average Shift (TAS) and Duration Average Shift (DAS) are defined analogously. Signedness is kept in order to identify the direction in which the underlying OMR system tends to move the notes. These should be accompanied by the corresponding standard deviation.
In order for all of these metrics to be independent of the sequence length, they should be computed and accumulated for the entire dataset and not averaged on a by-prediction basis.
There is also the possibility of extending the TER metric with some semantics in order to ensure better matching and more fine-grained information from it. Instead of matching the predicted and ground truth trees with their structure alone, it would be interesting to incorporate pitch into the notehead elements. Thus, a pitch change can be though of a partial substitution of the node at a reduced cost of $0.5$ for either the staff, the position on the staff or the notehead type. If more than one of this properties change, it should be considered a full substitution.
6 The COMREF Dataset
We have developed a dataset using our notation built on transcriptions of public domain works as a proof of concept of the notation format. In particular, we have used the OpenScore project’s transcriptions of widely known works such as The Art of the Fugue by J.S. Bach or the Planets by Gustav Holst, among others. We have also incorporated the Lieder Corpus Gotham and Jonas (2022) and the String Quartet corpus str (2023). All these scores are engraved in MusicXML, but we plan on supporting MEI as a conversion source at some point in the future.
6.1 Images
In summary, the dataset is produced by processing of 896 individual works into images at the measure level (including all staves that belong to it), to produce a total of $439.972$ images. After verification, $3.260$ images are removed due to noise in the annotations, totalling $436.752$ images after cleanup.
Figure 4: The pipeline through which the COMREF dataset has been generated.
The process through which the dataset was produced is summarised in Figure 4. Scores are engraved through Verovio into page-level SVG files. Using the hierarchical structure within the SVG and exploiting the optional identifier information Verovio can be instructed to attach, we employ a splitter script that finds bounding boxes for every measure and engraves them individually. It also finds what measures lie at the beginning of a line in order to properly add this information on the notation. The bounding boxes for all measures are cropped at the next staff vertically and a fixed amount of pixels horizontally.
Once the images are produced, the converter uses the MusicXML file and produces the MTN notation. In order to ensure all images have their corresponding ground truth, we use a cleaning script that finds matching identifiers for images in the MTN files. It also checks for outliers in case there are blatant mistakes in the notation. Although we have taken precautions to minimise the number of errors, there are a few images with objects far from the staff, either temporally or graphically. We remove these outliers heuristically to ensure the quality of the data.
6.2 Ground Truth
The dataset incorporates 91 different classes, which are shown alongside their frequencies in Table 3. A relevant aspect of this dataset is that it is quite unbalanced, with roughly 70% of the elements pertaining to the 4 first classes.
Table 3: Tokens and their frequencies in the COMREF dataset.
Token Count Token Count Token Count
notehead_black 1986948 rest_breve 18103 notehead_white_cue 738
beam 1517194 clef_C 17664 rest_whole 681
stem_down 1018045 notehead_black_grace 13086 regular 529
stem_up 803884 tenuto 11444 sfz 462
flag 244232 strong_accent 9414 rf 173
accidental_sharp 226428 pp 8484 mordent 136
accidental_flat 211192 rest_32nd 8381 ffff 130
notehead_white 197408 ff 7449 sfp 129
staccato 171070 sf 7282 notehead_breve 125
slur_stop 168855 light_light 7230 sffz 105
slur_start 168853 notehead_black_cue 6889 clef_G_oct1 102
dot 153128 rest_maxima 6171 turn 83
rest_eighth 126179 mf 5748 caesura 75
tied_start 101092 fermata 5330 slide_start 72
tied_stop 101029 light_heavy 3294 slide_stop 71
accidental_natural 99920 rest_64th 3248 accidental_natural_sharp 51
rest_quarter 94522 trill_mark 2589 rfz 43
clef_G 75140 timesig_common 2366 segno 31
clef_F 46386 accidental_double_sharp 1990 pppp 15
tuplet_stop 43781 fz 1700 glissando_stop 14
tuplet_start 43650 clef_G_oct-1 1295 glissando_start 10
rest_long 43610 staccatissimo 1268 clef_F_oct1 9
wedge_stop 38286 mp 1245 notehead_black_grace_cue 8
accent 27450 wavy_line 1107 coda 7
arpeggiate 22102 fff 987 schleifer 4
rest_half 21554 accidental_flat_flat 969 sfpp 3
rest_16th 21452 fp 913 accidental_natural_flat 2
p 21103 timesig_cut 890 pppppp 1
wedge_crescendo 20115 ppp 862 ppppp 1
f 19771 clef_percussion 774
wedge_diminuendo 18233 clef_F_oct-1 760 Total 8015354
7 Conclusions
In this paper we have argued for the implementation of a Common Optical Music Recognition Evaluation Framework through the instauration of a notation format in which score reconstruction can be evaluated fairly an unambiguously. Our proposed reification of this idea is the MTN format, which draws heavy inspiration from existing formats and adapts many ideas already present in the field while also being quite expressive and general. Its main idea is exploiting the expresiveness of graphs while keeping the ordering properties of sequences, thanks to which evaluation and inference are better-defined and less computationally demanding. Of course, CWMN is a tremendously complex notation system which has been optimised and streamlined for hundreds of years. It is highly likely that MTN (or any other computational representation of music for that matter) cannot model every possible music score that has been written. Nevertheless, we believe the subset of music that can be expressed in this format is large enough to be useful for the community of OMR.
In this work we have also presented a concrete implementation of a set of metrics for OMR practitioners with the hopes of bringing together the community to finally speak the same language; a lingua franca thanks to which research can be shared and compared fairly and easily. We also hope this might lower the barrier of entry and bring some new interest on the field.
The work that lies ahead now is bulding a corpus of music that can be employed with this format into a benchmark for CWMN recognition, both in typeset and handwritten domains. For typeset scores, building a corpus in MTN should be as straightforward as compiling a comprehensive set of scores in MusicXML and engraving them. For handwritten music, the million dollar question now lies in how to reliably align MTN to the source material and where to find good transcriptions, which shall be the focus of our next efforts.
[…]
Declarations
The authors have no conflicts of interest to declare that are relevant to the content of this article.